OpenStudy (anonymous):

What is the value of x in the following equation, 1/5x = 10? 2 5 15 50

OpenStudy (studygurl14):

is it \[\ (\frac{ 1 }{ 5 })x = 10\] or \[\frac{ 1 }{ 5x } = 10\]

OpenStudy (anonymous):

It's the first one but without parenthesis.

OpenStudy (studygurl14):

okay. So then you would do... \[\frac{ x }{ 5 } = 10\] \[\ 5(\frac{ x }{ 5 }) = 5(10)\] \[x = 50\]

OpenStudy (shamil98):

multiply both sides by 5 to get rid of the fraction and find x.

OpenStudy (anonymous):

Oh thanks @StudyGurl14 :)

OpenStudy (anonymous):

Hi @shamil98

OpenStudy (studygurl14):

you're welcome @shamil98

OpenStudy (studygurl14):

Sorry, I meant to put @Trojan

OpenStudy (anonymous):

lol :) @StudyGurl14